# Exploiting Minority Pseudo-Labels for Semi-Supervised Semantic Segmentation in Autonomous Driving

### v2

@@ 3.3 Dual Contrastive Learning with Anchors
Real-world datasets frequently display long-tail distributions, leading to significant class imbalances where dominant classes can distort model training and influence decision boundaries for minority classes. To address this issue, we propose a dual contrastive learning with anchors. This strategy aims to maintain a uniform distribution of all classes, including minority classes, within the feature space. By doing so, it enhances decision boundaries and promotes better generalization, particularly effective for handling long-tailed data distributions.
Predefined Anchors. First, we define $C$ as the number of classes and $v_{i}$ with $i=1,2,3,...,C$ as anchors. The dimension of each anchor is equal to the dimension of the features. These anchors are randomly initialized, and to ensure that each has the same distance as the others, we perform the following minimization approach: [42]:
$\mathcal{L}_{Anchor}=\frac{1}{C}\sum_{i=1}^{C}log\sum_{j=1}^{C}e^{v_{i}^{T}\cdot v_{j}/\tau},$ (12)
where $\tau$ is a temperature coefficient.
We also define $c_{i}^{t}$ as the prototype of class $i$ at time step $t$. To obtain the appropriate anchor for each class, we calculate the prototype of each class at the early stages of the training and iteratively update them by the EMA:
$c_{i}^{t}=\alpha(c_{i}^{t-1})+(1-\alpha)(c_{i}^{t}),$ (13)
where $\alpha$ is 0.99 by default to ensure that the prototypes were relatively stable. A prototype $c_{i}$ is matched with the anchor index $v_{\sigma_{i}}$ with the following constraints:
$\sigma=\arg\min_{\sigma}\sum_{i=1}^{C}\left\|v_{\sigma_{i}}-c_{i}\right\|,$ (14)
where $\sigma$ is a distributive permutation of vector $[0,1,2,...,N]$ that represents a one-to-one correspondence between the class and the anchor. This optimization can be solved by the Hungarian algorithm [43].
@@
Proximity-Guided Sampling. The memory bank stores representative features from labeled data and is used for contrastive learning [40, 25, 15]. Due to space limitations, the proposed memory bank only stores a subset of representative features obtained from labeled data, which comprises a first-in-first-out (FIFO) queue for each class. Based on this, we propose a proximity-guided sampling strategy to obtain a more compact memory bank. We extract features only from Gen-Teacher by a projection head. To select the subset of features to be included in the memory bank, we predefine a threshold $\phi$ to choose the features with higher confidence, and then those features closer to their corresponding anchors are preferred. We update the memory bank only with the $top-K$ closest features and pop out the outdated features at the top.
@@
$\mathcal{L}_{sim}=\frac{1}{\left|M_{c}\right|}\sum_{i^{+}\in Q_{c}}(1-\frac{\left\langle f,i^{+}\right\rangle}{\left\|f\right\|_{2}\cdot\left\|i^{+}\right\|_{2}}),$ (16)
where $f$ represents the feature of the predicting pixel belonging to class $c$, $i^{+}$ denote the representative features of class $c$ in the memory bank $Q_{c}$.
3.4 Training Process
@@
4 Experimental Results
In this section, we assess the effectiveness of our proposed method on two well-established datasets: the Cityscapes dataset [44] and the PASCAL VOC 2012 dataset [45]. The experiments are performed using PyTorch on a server equipped with an NVIDIA A40 GPU.
4.1 Experimental Setting
Datasets. The Cityscapes dataset [44], designed for semantic analysis of urban street scenes, contains 2975 finely annotated training images and 500 validation images across 19 classes. However, the dataset suffers from a significant long-tail problem, with classes like roads and sky occupying the majority of pixels, while others like pedestrians and poles account for much fewer pixels, which severely hinders the model’s learning performance. In addition, due to the limitations of open source datasets on autonomous driving, we also verified the effectiveness of our approach on PASCAL VOC 2012 dataset which [45] serves as a standard benchmark for semantic segmentation with 20 object classes and 1 background class. And it includes 1464 training images and 1449 validation images. Additionally, we use the SBD dataset [46] as an augmented set, which provides 9118 extra training images. Due to the coarse annotations in the SBD dataset, previous methods employ two primary partitioning strategies. The first strategy uses only the standard 1464 images as the entire labeled set, while the second strategy considers all 10582 images as potential labeled data. To ensure a fair comparison of our methods, we evaluate them using both a blender set (10582 potential labeled images) [15] and a classic set (1464 labeled images) [11].
@@ 4.1 Experimental Setting
Implementation Details. The two student models share the same architecture, utilizing DeepLabV3+, but have different initializations. Both models are trained using the stochastic gradient descent (SGD) optimizer. When training on the Cityscapes dataset, the learning rate is initialized to ${10^{-2}}$, whereas, for the PASCAL VOC 2012 dataset, it is $5\times 10^{-3}$. The momentum for the optimizer is maintained at $0.9$. For adjusting the learning rate, we employ a polynomial decay strategy defined by $1-(\frac{\text{iter}}{\text{max\_iter}})^{\text{power}}$, with the power parameter set to $0.9$. During training, images are randomly cropped to $800\times 800$ pixels with a batch size of 3 for the Cityscapes dataset. For the PASCAL VOC 2012 dataset, a crop size of $512\times 512$ pixels with a batch size of 8 is used. The weights and hyperparameters for the loss functions are configured as follows: $\lambda_{s}=1$, $\lambda_{u}=1$, $\lambda_{ctr}=0.1$, $\tau=0.5$, $\phi=0.95$, and $N=256$. To implement strong data augmentation, three rectangular regions with random ratios ranging from 0.25 to 0.5 are randomly positioned within the input image and augmented using the CutMix [48] strategy.
4.2 Results
@@
Results on the Cityscapes dataset. We evaluated the effectiveness of STPG using the Cityscapes dataset. To ensure fairness, DeepLabV3+ with ResNet50 is employed across all comparison methods. In all experiments, we arbitrarily choose 1/16, 1/8, and 1/4 of the Cityscapes training set as labeled data, corresponding to 186, 372, and 744 images respectively, with the remaining portion as unlabeled data. Table 1 presents the comparative results, where STPG demonstrates a substantial improvement over existing methods. Specifically, STPG enhances the mIoU by 8.03%, 8.40%, and 5.85% in the 1/16, 1/8, and 1/4 columns, respectively, compared to baseline methods using only labeled data.
@@
Table 1: Comparisons of our STPG with the SOTA methods on the Cityscapes validation set using various partition protocols. We arbitrarily choose subsets from the training set of Cityscapes to use as labeled data: 1/16, 1/8, and 1/4, corresponding to 186, 372, and 744 images respectively, while the remaining training data serves as unlabeled. Each method employs DeepLabV3+ with ResNet50.
Methods 1/16 (186) 1/8 (372) 1/4 (744)
Sup. baseline 60.43 66.74 71.16
CAC [49] - 69.7 72.7
PC2Seg [41] - 72.1 73.8
ELN [50] - 70.3 73.5
ST++ [51] - 72.7 73.8
PGCL [52] - 71.2 73.9
MT [17] 66.1 72.0 74.5
CCT [21] 66.4 72.5 75.7
GCT [16] 65.8 71.3 75.3
MMFA [39] 66.2 73.6 76.9
STPG(Ours) 68.46 75.14 77.01
Results on the blender PASCAL VOC 2012 Dataset. We further assessed the performance of STPG on the blender PASCAL VOC 2012 dataset, ensuring all comparison methods utilized DeepLabV3+ with ResNet50 for consistency. In all experiments, we arbitrarily choose 1/16, 1/8, 1/4, and 1/2 of the PASCAL VOC 2012 training set as labeled data, corresponding to 662, 1323, 2646, and 5291 images respectively, with the remaining images treated as unlabeled data. Table 2 presents the comparative results. STPG consistently achieves state-of-the-art results across all partition protocols, surpassing existing best results by 0.08%, 0.19%, 1.17%, and 1.19% for the respective data splits.
@@
Table 2: Comparisons of our STPG with the SOTA methods on the blender PASCAL VOC 2012 validation set using various partition protocols. We arbitrarily choose subsets from the PASCAL VOC 2012 training set to use as labeled data: 1/16, 1/8, 1/4, and 1/2, corresponding to 662, 1323, 2646, and 5291 images respectively, while the rest of the training set serves as unlabeled data. Each method employs DeepLabV3+ with ResNet50.
Methods 1/16 (662) 1/8 (1323) 1/4 (2646) 1/2 (5291)
Sup. baseline 61.72 66.83 71.52 73.96
CutMix-Seg [20] 68.9 70.7 72.5 74.5
ADS-SemiSeg [53] - 70.8 72.8 -
ELN [50] - 73.2 75.6 -
CPS [11] 72.0 73.7 74.9 76.2
ST++ [51] 72.6 74.4 75.4 -
U2PL [15] 72.0 75.1 76.2 -
CPCL [54] 71.7 73.7 74.6 75.3
PGCL [52] - 75.2 76.0 -
MMFA [39] 68.0 72.5 75.7 -
VCc [55] 73.8 75.3 75.8 -
STPG(Ours) 73.88 75.49 76.97 77.39
@@
Results on the classic PASCAL VOC 2012 Dataset. We evaluate STPG on the classic PASCAL VOC 2012 dataset by selecting labeled images from the original fine-grained annotated training set (1464 images total) and treating all images from the SBD as unlabeled data. For fairness, all comparative methods employed DeepLabV3+ with ResNet101. In all experiments, we arbitrarily choose 1/8, 1/4, 1/2, and the full set of the PASCAL VOC 2012 training data as labeled, corresponding to 183, 366, 732, and 1464 images respectively, with the remainder as unlabeled. Table 3 shows the comparison results. STPG consistently outperforms the supervised-only baseline, with improvements of +17.97%, +15.38%, +11.72%, and +8.33% for the 1/16, 1/8, 1/4, and 1/2 partition protocols, respectively.
@@
Table 3: Comparisons of our STPG with the SOTA methods on the classic PASCAL VOC 2012 validation set using different partition protocols. We arbitrarily choose subsets from the fine-grained annotated training set of PASCAL VOC 2012 to use as labeled data: 1/8, 1/4, 1/2, and the full set, corresponding to 183, 366, 732, and 1464 images respectively. The remaining training set, including all SBD images, serves as unlabeled data. Each method employs DeepLabV3+ with ResNet101.
Methods 1/8 (183) 1/4 (366) 1/2 (732) Full (1464)
Sup. baseline 54.32 62.31 67.36 72.12
CutMix-Seg [20] 63.5 69.5 73.7 76.5
PseudoSeg [56] 65.5 69.1 72.4 73.2
PC2Seg [41] 66.3 69.8 73.1 74.2
CPS [11] 67.4 71.7 75.9 -
CTT [25] 71.1 72.4 76.1 -
ST++ [51] 71.0 74.6 77.3 79.1
U2PL [15] 69.2 73.7 76.2 79.5
PS-MT [57] 69.6 76.6 78.4 80.0
FPL [58] 71.7 75.7 79.0 -
PRCL [38] 74.4 76.7 - 78.2
STPG(Ours) 72.29 77.69 79.08 80.45
Enhanced Performance for Tail Classes. Due to the significant long-tail problem in the Cityscapes dataset, we conducted additional experiments on some of the tail classes (such as Wall, Fence and Pole, etc.) and compared STPG with the baseline and FixMatch [24]. As shown in Fig. 4, STPG can significantly improve the performance of the model on tail classes, thereby enhancing overall performance.
Comparison of T-SNE Visualization. We compare the data feature distributions produced by the baseline and STPG after T-SNE visualization. In Fig. 5, the features have been well-aligned and formed clusters by STPG. The model?s decision boundary becomes clearer, leading to more accurate predictions.
@@
Figure 4: Enhanced Performance for Tail Classes. The mIoU(%) for the tail classes (such as Wall, Fence, Pole, etc.) was evaluated using the baseline, FixMatch, and STPG. The results are derived from the Cityscapes dataset using 1/8 of the labeled data, employing DeepLabV3+ with ResNet50.

### v3

@@ III-C Anchor Contrastive Learning
Real-world datasets frequently display long-tail distributions, leading to significant class imbalances where dominant classes can distort model training and influence decision boundaries for minority classes. To address this issue, as shown in Fig. 5, we propose an anchor contrastive learning strategy using evenly distributed anchors and optimal anchor-prototype matching to alleviate majority-class feature dominance and thus enhance class discrimination. This strategy aims to maintain a uniform distribution in the feature space for all classes, including minority classes. By doing so, it enhances decision boundaries and promotes better generalization, which is particularly effective for handling long-tailed data distributions.
Figure 5: Illustration of anchor contrastive learning. We generate anchors that are evenly distributed in the feature space and perform one-to-one matching with class prototypes. During training, we sample a large number of features near the anchors and combine them with the anchors for contrastive learning.
Evenly Distributed Anchors Generation. Before training, we define $C$ as the number of classes and $v_{i}$ with i=1,2,3,,Ci=1,2,3,...,C as anchors. The dimension of each anchor is equal to the dimension of the features, denoted as $D$, i.e., $v_{i}\in\mathbb{R}^{D}$. These anchors are randomly initialized. To avoid any potential bias for initial values and ensure that they are in a comparable scale with the features, they are sampled from a standard normal distribution, i.e., $v_{i}\sim\mathcal{N}(0,I_{D})$, and then normalized to have unit L2-norm.
To achieve an evenly distributed initial state, after random initialization, we apply an optimization step to ensure that each anchor maintains a similar distance from others. Let $V=\{v_{1},v_{2},\dots,v_{C}\}$ be the set of $C$ anchor vectors. We then perform the following objective minimization approach, adapted from [58]:
$\min_{V}\left(\frac{1}{C}\sum_{i=1}^{C}\log\sum_{j=1}^{C}e^{v_{i}^{T}\cdot v_{j}/\tau}\right),$ (14)
where $T$ is matrix transpose. This objective function, designed with a temperature coefficient $\tau$, works to push anchors apart from one another, thereby ensuring they adopt a broadly spherical arrangement on the unit hypersphere.
Prototypes Update. In the early stage of training, we define $c_{i}^{t}$ as the prototype of class $i$ at time step $t$. The initial prototypes $c_{i}^{0}$ for each class $i$ are randomly initialized. Similar to the anchors, they are sampled from a standard normal distribution, and subsequently normalized to have unit L2-norm. This ensures that prototypes start in a general manner without initial bias from specific data instances.
To obtain the appropriate representation for each class, we continuously update the prototypes iteratively by the Exponential Moving Average (EMA). At each training step $t$, after computing the features for a mini-batch, for each class $i$ present in that mini-batch, we first calculate $f_{i,\text{avg}}$ as the average of the feature vectors of all samples belonging to class $i$ within this current mini-batch. Then, the prototype $c_{i}^{t}$ is updated as:
$c_{i}^{t}=\alpha(c_{i}^{(t-1)})+(1-\alpha)(f_{i,\text{avg}}),$ (15)
where $\alpha$ is a momentum coefficient. The prototypes are typically normalized after each update to maintain consistency with the anchors and features, especially if cosine similarity is used as a distance metric in subsequent stages.
Anchor-Prototype Matching. A prototype $c_{i}$ is matched with the anchor $v_{\sigma_{i}}$ with the following constraints:
$\bm{\sigma}^{*}=\arg\min_{\bm{\sigma}\in S}\sum_{i=1}^{C}\left\|v_{\sigma_{i}}-c_{i}\right\|,$ (16)
where $S$ denotes the set of all permutation of $\{1,2,...,C\}$, and $\bm{\sigma}=[\sigma_{1},\sigma_{2},...,\sigma_{C}]^{\top}$ is a permutation vector that assigns the $\sigma_{i}$-th anchor $v_{\sigma_{i}}$ to the prototype $c_{i}$ of the $i$-th class. The objective is to minimize the total distance between each class prototype $c_{i}$ and its matched anchor $v_{\sigma_{i}}$. This optimization problem seeks to determine the optimal $\bm{\sigma}^{*}$. Such a problem can be modeled as an assignment problem and can be efficiently solved by the Hungarian algorithm [33].
@@
Proximity-Guided Sampling. The memory bank stores representative features from labeled data and is used for contrastive learning [1, 62, 60]. Due to space limitations, the proposed memory bank only stores a subset of representative features obtained from labeled data, which comprises a first-in-first-out (FIFO) queue for each class. Based on this, we propose a proximity-guided sampling strategy to obtain a more compact memory bank. We extract features only from Gen-Teacher by a projection head. To select the subset of features to be included in the memory bank, we predefine a threshold $\phi$ to choose the features with higher confidence, and then those features closer to their corresponding anchors are preferred. We update the memory bank only with the topKtop-K closest features and pop out the outdated features at the top.
@@
$\mathcal{L}_{sim}=\frac{1}{\left|\mathcal{Q}_{c}\right|}\sum_{i^{+}\in\mathcal{Q}_{c}}(1-\frac{\left\langle f,i^{+}\right\rangle}{\left\|f\right\|_{2}\cdot\left\|i^{+}\right\|_{2}}),$ (18)
where $f$ represents the feature of the predicting pixel belonging to class $c$, $i^{+}$ denote the representative features of class $c$ in the memory bank $\mathcal{Q}_{c}$.
Complexity Analysis. We conduct time and space complexity analysis of dual contrastive learning. The dual contrastive objective consists of the anchor contrastive loss $\mathcal{L}_{ac}$ and the similarity loss $\mathcal{L}_{sim}$. For $\mathcal{L}_{ac}$, each feature vector $f$ interacts with one positive anchor and C1C-1 negative anchors. With feature dimension $D$, this yields a per-feature cost of $O(CD)$, and processing $N_{f}$ features per batch results in a total cost of $O(N_{f}CD)$. The memory required to store the class anchors is $O(CD)$. For $\mathcal{L}_{sim}$, each feature compares with $|\mathcal{Q}_{c}|$ stored samples of its class in the memory bank. Each cosine similarity computation costs $O(D)$, giving a per-feature cost of $O(|\mathcal{Q}_{c}|D)$ and a per-batch cost of $O(N_{f}|\mathcal{Q}_{c}|D)$. The memory footprint of the class-specific memory bank is $O(|\mathcal{Q}_{c}|D)$.
III-D Training Process
@@
IV Experimental Results
In this section, we evaluate the effectiveness of our proposed method on two widely-used semantic segmentation datasets for autonomous driving: the Cityscapes dataset [14] and the CamVid dataset [4]. In addition, to compare our approach with state-of-the-art methods, we also conduct performance evaluations on the PASCAL VOC 2012 dataset. The experiments are implemented using PyTorch on a server with an NVIDIA A40 GPU.
IV-A Experimental Setting
Datasets. There are two datasets for semantic segmentation in urban driving scenarios: The Cityscapes dataset [14], designed for semantic analysis of urban street scenes, contains 2975 finely annotated training images and 500 validation images across 19 classes, while the CamVid dataset [4], captured from a driving automobile’s perspective, includes 367 training and 101 validation images across 11 classes; however, both datasets exhibit a significant long-tail distribution, where dominant classes like roads and sky occupy the majority of pixels, whereas underrepresented classes such as pedestrians and poles comprise much fewer pixels, severely hindering model learning performance. In addition, to compare our approach with state-of-the-art methods, we also verified the effectiveness of our approach on PASCAL VOC 2012 dataset which [18] serves as a standard benchmark for semantic segmentation with 20 object classes and 1 background class. And it includes 1464 training images and 1449 validation images. Additionally, we use the SBD dataset [25] as an augmented set, which provides 9118 extra training images. Due to the coarse annotations in the SBD dataset, previous methods employ two primary partitioning strategies. The first strategy uses only the standard 1464 images as the entire labeled set, while the second strategy considers all 10582 images as potential labeled data. To ensure a fair comparison of our methods, we evaluate them using both a blender set (10582 potential labeled images) [60] and a classic set (1464 labeled images) [11].
@@ IV-A Experimental Setting
Implementation Details. Prior to the main training phase, we pre-optimize for 50 epochs using an Adam optimizer and a $10^{-3}$ learning rate to generate evenly distributed anchors. The two student models share the same architecture, utilizing DeepLabV3+, but have different initializations. Both models are trained using the stochastic gradient descent (SGD) optimizer. When training on the Cityscapes and CamVid dataset, the learning rate is initialized to $10^{-2}$}, whereas, for the PASCAL VOC 2012 dataset, it is $5\times 10^{-3}$. The momentum for the optimizer is maintained at $0.9$. For adjusting the learning rate, we employ a polynomial decay strategy defined by $1-(\frac{\text{iter}}{\text{max\_iter}})^{\text{power}}$, with the power parameter set to $0.9$. During training, images are randomly cropped to $800\times 800$ pixels with a batch size of 4 for the Cityscapes dataset. For the PASCAL VOC 2012 dataset, a crop size of $512\times 512$ pixels with a batch size of 8 is used. In the case of the CamVid dataset, the configuration remains similar to that of Cityscapes, except that the images are cropped to $360\times 480$ pixels and the batch size is set to 16. The weights and hyperparameters for the loss functions are configured as follows: $\lambda_{s}=1$, $\lambda_{u}=1$, $\lambda_{ctr}=0.1$, $\tau=0.5$, $\phi=0.95$, and $N=256$. To implement strong data augmentation, three rectangular regions with random ratios ranging from 0.25 to 0.5 are randomly positioned within the input image and augmented using the CutMix [71] strategy.
IV-B Results
@@
Results on the Cityscapes dataset. We evaluated the effectiveness of STPG using the Cityscapes dataset. To ensure fairness, DeepLabV3+ with ResNet50 is employed across all comparison methods. In all experiments, we arbitrarily choose 1/8, 1/4, and 1/2 of the Cityscapes training set as labeled data, corresponding to 372, 744, and 1488 images respectively, with the remaining portion as unlabeled data. Table I presents the comparative results, where STPG demonstrates a substantial improvement over existing methods. Specifically, STPG enhances the mIoU by 8.40%, 5.85% and 4.12% in the 1/8, 1/4, and 1/2 columns, respectively, compared to baseline methods using only labeled data.
@@
TABLE I: Comparisons of STPG with the state-of-the-art methods on the Cityscapes validation set using various partition protocols. We arbitrarily choose subsets from the training set of Cityscapes to use as labeled data: 1/8, 1/4 and 1/2, corresponding to 372, 744, and 1488 images respectively, while the remaining training data serves as unlabeled. Each method employs DeepLabV3+ with ResNet50. * indicates a reproduction under the same settings as ours.
Methods Year 1/8 (372) 1/4 (744) 1/2 (1488)
Sup. baseline - 66.74 71.16 74.36
MT [54] 2017 72.0 74.5 -
CCT [50] 2020 72.5 75.7 -
GCT [34] 2020 71.3 75.3 -
CAC [39] 2021 69.7 72.7 -
PC2Seg [75] 2021 72.1 73.8 -
ELN [38] 2022 70.3 73.5 75.3
ST++ [66] 2022 72.7 73.8 -
PGCL [37] 2023 71.2 73.9 76.8
MMFA [70] 2023 73.6 76.9 -
CPCL [19] 2023 74.6 76.9 -
UniMatch [65] 2023 75.7 76.8 78.1
MGCT [29] 2024 73.4 74.0 -
CorrMatch [53] 2024 74.1 76.8 77.1
UCCL [69] 2025 71.2 73.1 74.3
ScaleMatch [46] 2025 75.1 76.5 76.9
NRCR [72] 2025 73.5 73.9 75.8
MCCL [68] 2025 71.0 70.6 72.7
STPG(Ours) 2025 75.14 77.01 78.48
Results on the CamVid dataset. We evaluated the effectiveness of STPG using the CamVid dataset. To ensure fairness, DeepLabV3+ with ResNet50 is employed. In all experiments, we arbitrarily choose 1/8, 1/4, and 1/2 of the training set as labeled data, corresponding to 45, 91, and 183 images respectively, with the remaining portion as unlabeled data. Figure 6 presents the comparative results, where STPG demonstrates a substantial improvement over existing methods. Specifically, STPG enhances the mIoU by 5.40%, 4.96%, and 3.43% in the 1/8, 1/4, and 1/2 columns, respectively, compared to baseline methods using only labeled data.
@@
Figure 6: Performance comparison on the CamVid dataset. Comparisons of STPG with baseline and FixMatch on the CamVid validation set using various partition protocols. We arbitrarily choose subsets from the CamVid training set to use as labeled data: 1/8, 1/4, and 1/2, corresponding to 45, 91, and 183 images respectively, while the rest of the training set serves as unlabeled data. Each method employs DeepLabV3+ with ResNet50.
Results on the blender PASCAL VOC 2012 Dataset. We further assessed the performance of STPG on the blender PASCAL VOC 2012 dataset, ensuring all comparison methods utilized DeepLabV3+ with ResNet50 for consistency. In all experiments, we arbitrarily choose 1/16, 1/8, 1/4, and 1/2 of the PASCAL VOC 2012 training set as labeled data, corresponding to 662, 1323, 2646, and 5291 images respectively, with the remaining images treated as unlabeled data. Table II presents the comparative results. STPG consistently achieves state-of-the-art results across all partition protocols, surpassing existing best results by 0.08%, 0.19%, 1.17%, and 1.19% for the respective data splits.
@@
Results on the classic PASCAL VOC 2012 Dataset. We evaluate STPG on the classic PASCAL VOC 2012 dataset by selecting labeled images from the original fine-grained annotated training set (1464 images total) and treating all images from the SBD as unlabeled data. For fairness, all comparative methods employed DeepLabV3+ with ResNet101. In all experiments, we arbitrarily choose 1/8, 1/4, 1/2, and the full set of the PASCAL VOC 2012 training data as labeled, corresponding to 183, 366, 732, and 1464 images respectively, with the remainder as unlabeled. Table III shows the comparison results. STPG consistently outperforms the supervised-only baseline, with improvements of +17.97%, +15.38%, +11.72%, and +8.33% for the 1/16, 1/8, 1/4, and 1/2 partition protocols, respectively.
@@
TABLE II: Comparisons of STPG with the state-of-the-art methods on the blender PASCAL VOC 2012 validation set using various partition protocols. We arbitrarily choose subsets from the PASCAL VOC 2012 training set to use as labeled data: 1/16, 1/8, 1/4, and 1/2, corresponding to 662, 1323, 2646, and 5291 images respectively, while the rest of the training set serves as unlabeled data. Each method employs DeepLabV3+ with ResNet50. * indicates a reproduction under the same settings as ours.
Methods Year 1/16 (662) 1/8 (1323) 1/4 (2646) 1/2 (5291)
Sup. baseline - 61.72 66.83 71.52 73.96
CutMix-Seg [23] 2019 68.9 70.7 72.5 74.5
CPS [11] 2021 72.0 73.7 74.9 76.2
ADS-SemiSeg [5] 2022 - 70.8 72.8 -
ELN [38] 2022 - 73.2 75.6 -
ST++ [66] 2022 72.6 74.4 75.4 -
U2PL [60] 2022 72.0 75.1 76.2 -
CPCL [19] 2023 71.7 73.7 74.6 75.3
PGCL [37] 2023 - 75.2 76.0 -
MMFA [70] 2023 68.0 72.5 75.7 -
UniMatch [65] 2023 73.4 75.1 76.1 76.9
VC3 [27] 2024 73.8 75.3 75.8 -
MGCT [29] 2024 73.4 75.6 75.9 -
STPG(Ours) 2025 73.88 75.49 76.97 77.39
TABLE III: Comparisons of our STPG with the state-of-the-art methods on the classic PASCAL VOC 2012 validation set using different partition protocols. We arbitrarily choose subsets from the fine-grained annotated training set of PASCAL VOC 2012 to use as labeled data: 1/8, 1/4, 1/2, and the full set, corresponding to 183, 366, 732, and 1464 images respectively. The remaining training set, including all SBD images, serves as unlabeled data. Each method employs DeepLabV3+ with ResNet101. * indicates a reproduction under the same settings as ours.
Methods Year 1/8 (183) 1/4 (366) 1/2 (732) Full (1464)
Sup. baseline - 54.32 62.31 67.36 72.12
CutMix-Seg [23] 2019 63.5 69.5 73.7 76.5
PseudoSeg [78] 2020 65.5 69.1 72.4 73.2
PC2Seg [75] 2021 66.3 69.8 73.1 74.2
CPS [11] 2021 67.4 71.7 75.9 -
CTT [62] 2022 71.1 72.4 76.1 -
ST++ [66] 2022 71.0 74.6 77.3 79.1
U2PL [60] 2022 69.2 73.7 76.2 79.5
PS-MT [44] 2022 69.6 76.6 78.4 80.0
FPL [51] 2023 71.7 75.7 79.0 -
PRCL [64] 2023 74.4 76.7 - 78.2
SemiCVT [31] 2023 71.3 75.0 78.5 80.3
UPC [20] 2023 73.5 76.1 78.0 80.2
CSS [56] 2023 71.9 74.9 77.6 -
UniMatch [65] 2023 74.1 77.4 79.0 80.1
MGCT [29] 2024 73.8 77.1 78.2 80.3
STPG(Ours) 2025 72.29 77.69 79.08 80.45
Enhanced Performance for Tail Classes. Due to the significant long-tail problem in the Cityscapes dataset, we conducted an additional evaluation on some of the classes, especially the minority classes (such as Wall, Fence, Pole, etc.), and compared STPG with the baseline and FixMatch. As shown in Fig. 7, STPG can significantly improve the performance of the model on minority classes. Furthermore, by enhancing the learning of minority classes and introducing anchor-based contrastive learning, our framework enables the model to achieve better separation between minority and majority classes in the feature space. This improved feature separation not only benefits minority classes but also helps majority class samples to cluster more tightly in their respective regions, leading to marginal performance gains for majority classes as well. This observation highlights the strength of our framework: it effectively enhances the recognition performance of minority classes without compromising the accuracy of majority classes.
Comparison of T-SNE Visualization. We compare the data feature distributions produced by the baseline and STPG after T-SNE visualization. In Fig. 8, the T-SNE visualization of the model output features reveals that, compared to the baseline where feature distributions are scattered and class boundaries are blurred, our method demonstrates superior performance in feature alignment and clustering, with data points forming distinct class clusters and achieving better separation between classes; this observation indicates that our method can more effectively learn discriminative feature representations, leading to more accurate predictions.
@@
Figure 7: Enhanced performance for different classes. The mIoU(%) for the tail classes (such as Wall, Fence, Pole, etc.) and majority classes (such as Building, Sky, and Car), was evaluated using the baseline, FixMatch, and STPG. The results are derived from the Cityscapes dataset using 1/8 of the labeled data, employing DeepLabV3+ with ResNet50.